40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m elog  Not logged in ELOG logo
Entry  Mon Jul 20 11:57:17 2015, ericq, Update, General, Arm Locking recovered 
    Reply  Wed Jul 22 15:09:40 2015, ericq, Update, General, IMC work 
       Reply  Mon Aug 10 11:57:45 2015, ericq, Update, General, IMC work 
Message ID: 11436     Entry time: Wed Jul 22 15:09:40 2015     In reply to: 11430     Reply to this: 11486
Author: ericq 
Type: Update 
Category: General 
Subject: IMC work 

The other night, I spent some time with the mode cleaner. 

First, I made some model changes to the MC_TRANS part of c1mcs.mdl. Specifically, I brought in the userapps QPD part that we are using for the transmon and oplev QPDs. My main motivation for doing so was to have FMs for the pitch and yaw values, to be able to set offsets. Up until now, we have used a QPD centering servo in conjunction with the WFS, but the center of the QPD is not perfectly aligned to represent the center of MC2. Using offsets at the servo isn't really practical, since there are integrators. 

I then spent some time manually aligning the mode cleaner mirrors with WFS off, followed by centering the in-lock MC REFL beam on the WFS QPDs, and setting the WFS and MC_TRANS offsets. (I updated the WFS offset script, and made one for MC_TRANS in scripts/MC/WFS. They now use averaging instead of servoing to zero, a la LSC PD offset script). 

The resultant intracavity power and RIN was an improvement over the older offsets. (RMS RIN went down by half, I think.)

Since Monday, the autolocker seems to be having some trouble. At first, I suspected the changes I made weren't so hot after all, but I've now noticed a pattern. Often when I come to manually lock the mode cleaner due to a long unlocked period, I find that the sliders are not in the state specified by the mcdown script. Furthermore, it's not the same channels every time; sometimes the servo gain is left high, sometimes the boosts are left on. I fear that some of the caput commands are failing to execute. Ugh. 

ELOG V3.1.3-